Snoopy flies his Sopwith Camel the 120 kilometers from London to Calais in 12 hours, flying into a headwind. The return trip, with the same wind as a tailwind, takes only 2 hours. How fast is Snoopy flying throught the air? How fast is the wind blowing?
I would really like to see the solution and answer to this problem!

Found 2 solutions by Paul, checkley71:
Answer by Paul(988)

Let the speed of the Airplane be x
Let the speed of the wind be y
12(x-y)=120 (first equation)
2(x+y)=120 (second equation)
Solve for y in the 2nd equation:
2x+2y=120
2y=120-2x
y=60-x (subsitution)

From first equation:
12x-12y=120
Subsitute for y:
12x-12(60-x)=120
12x-720+12x=120
24x=840
x=35
y=60-35
y=25

Hence, he is flying at rate of 35mph with wind blowing at 25mph.
Paul.

Answer by checkley71(8403)

120/12=10 kph into the head wind
120/2=60 kph with the tail wing
therefore the wind is 25kph or 10+25=60-25 or 35=35 or 35 kph is the speed with no wind.